1. ysin 2x dx-dy-y^2dy-cos^2xdy=0
1
Expert's answer
2021-05-10T13:08:13-0400

The given differential equation is


"y sin2x dx -(1+y^2+cos^2x)dy=0"

Which is of the form "Mdx +Ndy =0" .

Where 


"M=y sin2x \ \text{and} \ N=-(1+y^2+cos2x)"



"\frac{\partial{M}}{\partial{y}}=\frac{\partial{ysin2x}}{\partial{y}}=sin2x""\frac{\partial{N}}{\partial{x}}=\frac{\partial{-(1+y^2+cos^2x)}}{\partial{x}}=-(2cosx×-sinx)"

"=2sinxcos =sin2x"x


As "\frac{\partial{M}}{\partial{y}}=\frac{\partial{N}}{\partial{x}}" ,Hence the given differential equation is an exact differential equation.


"\therefore \" The solution of the given differential equation is


"\int_{y=const.}Mdx+\int \text{(only those therms of N which do not contains x)}dy"

"=\int_{y=const.}ysin2x+\int{-(1+y^2)}dy"


"=y\int sin2x dx-\int(1+y^2)dy"

"=-\frac{ycos2x}{2}-y-\frac{y^3}{3}+c"

Where "c" is the integration constant